Opened 6 years ago

Closed 3 years ago

#452 closed flight processing (fixed)

BGS11/01, flight day 236b/2012, East Greenland

Reported by: knpa Owned by:
Priority: alpha 4 high Milestone: 2012 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description (last modified by anhi)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/greenland_iceland/BGS11_01-2012_236b_East_Greenland/

Data arrived from ARSF via SATA disk on 24/09/2012

BGS11/01 Scientific objective: unknown

BGS11/01 Priority: ?

PI: Tapsa Tukiainer

Sensors:

Camera (04/01/2013)
Eagle (27/03/2013)
Hawk (22/02/2013)
LiDAR (04/01/2013)

Change History (32)

comment:1 Changed 6 years ago by knpa

  • Description modified (diff)
  • Priority changed from alpha 4 medium to alpha 4 high

comment:2 Changed 6 years ago by edfi

finished navigation processing, created sol files incl. camera. Couldn't get good basestation so used PPP

comment:3 Changed 6 years ago by knpa

Berin processing HS.

comment:4 Changed 6 years ago by knpa

  • Description modified (diff)

comment:5 Changed 6 years ago by besm

Started processing LiDAR (required for HS). Used PPP navigation and 206a boresight values. Result is surprisingly good, so using the pitch+roll error from 206a, and not pursuing navigation data any further.

comment:6 Changed 6 years ago by besm

Actually, it looks as though pitch error at least results in a 1m+ offset on the ground. Returning to alspp to apply correction.

comment:7 Changed 6 years ago by besm

Decided on these pitch and roll errors.

Pitch: -0.003366
Roll: +0.0006162

They seem unusually low and high, respectively, but I'm attributing that to the fact that this data was processed using PPP.

comment:8 Changed 6 years ago by besm

Found other places in the data where the previous (outlying) results were radically wrong. I am compromising halfway between this and the 206a boresight, which appears to produce excellent results. There may have been an error in the calculation of the previous pitch and roll errors. The ones I'm finally settling on (for sure this time), are...

Pitch: -0.00317
Roll: +0.000423

comment:9 Changed 6 years ago by anhi

RCD Delivery Created

comment:10 Changed 6 years ago by besm

LiDAR delivery created.

comment:11 Changed 6 years ago by anhi

LiDAR Delivery check:
Hard to see LiDAR data in dem but can see the edge of the flight lines over the water so it is there.

No other problems found

comment:12 Changed 6 years ago by edfi

Beginning RCD delivery check

comment:13 Changed 6 years ago by besm

SCTs finally completed. Had a lot of trouble processing eagle line 5 using LiDAR ASTER DEM.

comment:14 Changed 6 years ago by edfi

LiDAR and Camera on disk 146 to be sent with other data for same project

comment:15 Changed 6 years ago by besm

Hyperspectral delivery finally finished and ready for delivery checking, after the umpteenth processing attempt.

comment:16 Changed 6 years ago by anhi

Hyperspectral Delivery check:

  • Moved logsheet in.
  • Hawk line 2 has dropped scan.
  • Eagle line 2 has a part dropped scan.
  • Eagle lines 5 and 7 xml file dem information is incorrect, proj_tidy found 'None' in them.

comment:17 Changed 6 years ago by edfi

  • Description modified (diff)

comment:18 Changed 6 years ago by besm

Ammended noted issues with hyperspectral data. Ready to re-check.

comment:19 Changed 6 years ago by mark1

  • Description modified (diff)

comment:20 Changed 6 years ago by anhi

Hyperspectral Delivery check:

  • No problems found, ready for delivery.

comment:21 Changed 6 years ago by besm

Remade a clean delivery as discussed to assert that the data is consistent with the documentation. Latest delivery is ready to check.

comment:22 Changed 6 years ago by anhi

Hyperspectral Delivery check:

  • hyp_genreadme-airborne.cfg needed updating for the new delivery, the new readme may be using old screenshots.

No other problems.

comment:23 Changed 6 years ago by besm

Readme is updated for hyperspectral delivery. Looks ok at a glance.

comment:24 Changed 6 years ago by edfi

Readme appears to all be fine. Ready to deliver

comment:25 Changed 6 years ago by edfi

Mapped files currently zipping

comment:26 Changed 6 years ago by edfi

Mapped files zipping complete

comment:27 Changed 6 years ago by anhi

Delivered Hawk data on 22/02/2013 on disk 151 to Tapani Tukiainen

comment:28 Changed 6 years ago by anhi

  • Description modified (diff)

comment:29 Changed 6 years ago by anhi

Delivered Eagle data on 27/03/2013 on disk 155 to Tapani Tukiainen

comment:30 Changed 6 years ago by anhi

  • Description modified (diff)

comment:31 Changed 4 years ago by tec

Archiving
Started

Ready for upload

Last edited 4 years ago by tec (previous) (diff)

comment:32 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Archived data now available at NEODC

Note: See TracTickets for help on using tickets.